The point P(21, 28) is on the terminal side of θ. Evaluate sin θ.
1
Expert's answer
2019-07-17T11:33:42-0400

Note that for P(21, 28) we get x=21 and y=28.

Consider O(0; 0), then "OP=\sqrt{21^2+28^2}=" "\sqrt{441+784}=\sqrt{1225}=35".

sin θ="\frac{opposite side}{hypotenuse}=\frac{y}{OP}=\frac{28}{35}=\frac{4}{5}."

Answer: sin θ"=\frac{4}{5}" .